OpenStudy (anonymous):

3/4 - 1/6 + 9/12=?. expressing final answer in reduced to lowest terms. please show all work.

OpenStudy (anonymous):

3/4 -1/6+ 9/12 = 9/12 -2/12 + 9/12 =16/12= 4/3